You have 96 feet of fencing to enclose a rectangular pen for your dog. To provide sufficient running space for your dog to exercise the pen should be 3 times long as its wide find the dimensions of the pen

Answer by rfer(16322)

2x+2(3x)=96
2x+6x=96
x=96/8
x=12
3x=36